Session: Feb 2000 (session 1)
Wavelength: 3.6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (A.Scalambra at Fri Feb 25 10:10:40 MET 2000) Success
sky clear

Onsala (A. Polatidis at Fri Feb 25 11:28:23 MET 2000) Success
Snowfall during the last 2 hours, otherwise clear skies. We can only record RCP and hence all VCs contain RCP.

Effelsberg (A. Kraus at Fri Feb 25 16:30:21 MET 2000) Minor failures
Due to a problem with the tape drive the first run is probably lost. From UT0620-UT0750 the telescope had to be stopped (strong winds).

OAN\-Yebes (Paco Colomer at Mon Feb 28 10:09:23 MET 2000) Failure
OAN-Yebes, in principle scheduled for this experiment, was not included in the drudg file. A warning sent to the EVN scheduler was not replied, and the experiment was not observed.

Westerbork (Tony Foley at Tue Feb 29 13:47:30 MET 2000) Minor failures
Several serious problems restarting the tied array after tests with new software., but started on time. Phasing of array maybe poor.

Westerbork (Tony Foley at Tue Feb 29 13:58:28 MET 2000) Minor failures
All scans between 05:15 and 08:37 out of hour angle range. After 08:39 all scans out of hour angle range. 8MHz USB and 8MHz LSB come from different IFs of the tied array, and so should not be combined before phase corrections.

Noto (Noto VLBI friend at Thu Mar 2 11:53:19 MET 2000) Success

Calibration (Alastair Gunn at Wed Apr 19 13:06:58 METDST 2000) Success
Calibration data available from vlbeer: see file feb00/gb034.antab for details